OpenStudy (anonymous):

Type the ordered pair that is the solution to these equations. 2x + 5y = -9 3x - 4y = -2

OpenStudy (anonymous):

2x=-9-5y x=(-9/2)-(5/2)y 3((-9/2)-(5/2)y)-4y=-2 (-27/2)y-(15/2)y-4y=-2 -25y=-2 y=2/25 x=-47/10

OpenStudy (anonymous):

so what would the ordered pair be?

OpenStudy (anonymous):